OpenStudy (anonymous):

The slope of any line perpendicular to , where A(0, -3) and B(4, 9) are two points on the line, is m =

OpenStudy (anonymous):

y-y1)/x - x1 = m

OpenStudy (anonymous):

but since you want a perpendicular line you use ., m1.m2 = -1

OpenStudy (anonymous):

The slope of the perpendicular line is the negative reciprocal of the slope of the original line, so, do let n be the slope of the line containing points A and B , then n = (9 - (-3))/(4 - 0) = 3, so m = -1/3